OpenStudy (anonymous):

i need step by step instrustions on what to do when asked to find the exact solutions of the equation in the interval 0,2pi when the equation is 4sin x cos x

OpenStudy (anonymous):

4sinx cosx =1

OpenStudy (anonymous):

It's a tougher problem...but here's what you do:

OpenStudy (anonymous):

i need to figure out how to do these for my test tomorrow

OpenStudy (anonymous):

sin (2x) = 2 sin x cos x (that's a double-angle formula) So, 4 sin x cos x is just 2 sin(2x) because 2 sin(2x) = 2(2 sin x cos x) = 4 sin x cos x So rewrite the equation as 2 sin(2x) = 1 Now, divide both sides by 2, you get sin(2x) = 1/2 The question is..the sine of what angle is 1/2... Answer many angles, like pi/6, 5pi/6, etc.. So take the angle 2x abd set is to pi/6

OpenStudy (anonymous):

2x = pi/6 x = pi/12...so pi/12 is one solution. Now, set 2x = 5pi/6 so solving for x, you get x = 5pi/12 so 5pi/12 is another solution to the problem.

OpenStudy (anonymous):

continue this process till you find all solutions in (0, 2pi).

OpenStudy (anonymous):

ahh ok. i see

OpenStudy (anonymous):

that wopuld be all solutions if i believ so right?

OpenStudy (anonymous):

You had to realize that 4 sin x cos x = 2 sin(2x)

OpenStudy (anonymous):

No, there are more solutions than pi/12 and 5pi/12. You will find the remaining solutions...

OpenStudy (anonymous):

i got to 2(2 sinx cosx) but didnt know what to do after that

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

7pi over 12 and 11pi/12 is the last two right?

OpenStudy (anonymous):

sekar is this right?